the length of a rectangle is 7 meters more than 4 times its width. the perimeter is 146 meters. find the dimensions of the rectangle.
Answer by Maths68(1474)

Length = l=4w+7 meters
Width = w=?
Perimeter = p = 146 meters

Perimeter of a Rectangle = length + length + width + width
Perimeter of a Rectangle =2 length +2 width
P = 2l +2w
146=2(4w+7)+2w
146=8w+14+2w
146=10w+14
146-14=10w
132=10w
132/10=10w/10
13.2=w
w=13.2 meters


Dimension of the rectangle.
Length = l=4w+7 = 4(13.2)+7 = 52.8+7 = 59.8 meters
Width = w=13.2 meters
Perimeter = p = 146 meters